# A Generic Machine Learning Framework for Radio Frequency Fingerprinting

### v1

@@ Abstract
Fingerprinting Radio Frequency (RF) emitters typically involves finding unique emitter characteristics that are featured in their transmitted signals. These fingerprints are nuanced but sufficiently detailed, motivating the pursuit of methods that can successfully extract them. The most granular downstream task is known as Specific Emitter Identification (SEI), which requires a well informed RF fingerprinting (RFF) approach for it to be successful. RFF and SEI have a long history, with numerous application areas in defence and civilian contexts such as signal intelligence, electronic surveillance, physical-layer authentication of wireless communication devices, to name a few. RFF methods also support many other downstream tasks such as Emitter Data Association (EDA) and RF Emitter Clustering (RFEC) and are applicable to a range of transmission types. In recent years, data-driven approaches have become popular in the RFF domain due to their ability to automatically learn intricate fingerprints from raw data. These methods generally deliver superior performance when compared to traditional techniques. The more traditional approaches are often labour-intensive, inflexible and only applicable to a particular emitter type or transmission scheme. Therefore, we consider data-driven Machine Learning (ML)-enabled RFF. In particular, we propose a generic framework for ML-enabled RFF which is inclusive of several popular downstream tasks such as SEI, EDA and RFEC. Each task is formulated as a RF fingerprint-dependent task within the framework. A variety of use cases using real RF datasets are presented here to demonstrate the framework for a range of tasks and application areas, such as spaceborne surveillance, signal intelligence and countering drones.
Applied Intelligence Labs, ITG, Defence
             Digital Intelligence, BAE Systems, Chelmsford, Essex, CM2 8HN
1 Introduction
@@
RFF involves building emitter representations from characteristics that are present in their signals. This can be particularly useful for downstream tasks such as SEI, which involves classifying signals into their emitter origination (e.g. differentiating between various radars, 5G sources, etc) [21, 44]. In general, the emitter characteristics are unintentional (e.g., the waveform, transmission timing or bandwidth), caused by the channel (e.g., Doppler frequency offset, dispersion, scintillation), or intentional [21]. Intentional characteristics are often assumed to originate from inherent transmitter hardware imperfections (e.g., amplitude rise-time, ripple, phase distortion, nonlinearities in the transmitter processing chain, antenna responses) and/or are induced by manufacturing variabilities, including individual components in the transmitter processing chain. Spoofing or concealing these characteristics is difficult, allowing tasks such as SEI to be possible. Overall, RFF involves finding set of specific characteristics or features (intentional or not) which are unique to a RF emitter that can be extracted from the received transmission [21, 44].
In practice, datasets containing received transmissions are collected, and subsequently used by a RFF approach in order to extract emitter characteristics. These datasets also contain different types of transmissions or modulation schemes, allowing us to solve other less relevant identification tasks such as distinguishing between types of transmissions or types of modulation. The primary goal here, however, is to obtain RF fingerprint representations which allow us to solve tasks such as SEI, since solving this task requires a detailed account of the intentional features originating from the emitters’ hardware imperfections. More generally, emitter classification can be on the individual hardware unit level (e.g. SEI) or even coarser (e.g. individual manufacturers). Consequently, RFF for SEI is the most detailed and nuanced, typically exploiting unintentional features originating from the emitters’ hardware imperfections and/or manufacturing variabilities to solve the SEI task successfully. Note that broader emitter class characteristics such as type, modulation scheme, central frequency, used waveform, etc may also be used either implicitly or explicitly to inform RFF.
In general, there are two different ways to approach RFF:
@@
    Traditional RFF with signal processing techniques [4, 19, 9, 25, 13, 8],
@@
    Deep Learning (DL)-based, ML, RFF [47, 45, 42, 50, 15].
@@
The first approach manually defines unique features/characteristics that form the RF fingerprint for a sought set of emitters (e.g., evolution of the signal’s frequency spectrum or its amplitude envelope). Subsequently, a suitable conventional signal processing (SP) algorithm is then applied to extract this RF fingerprint. In general, such conventional approaches to RFF can be effective and their outputs fully explainable. However, these techniques are inherently inflexible and are usually specific to a particular type of emitter/signal or classification requirement. In addition, they require considerable effort by specialists for each application or emitter/signal type. For example, they may need to manually define the new RF fingerprint and devise or adapt the associated SP method. This renders their development cycle for new types or classes of emitters slow, labour-intensive and expensive. This is particularly problematic in the increasingly congested and contested electromagnetic environment (EME), where new emitters emerge regularly and communication technologies continuously evolve. Therefore, a more versatile and flexible approach to RFF might be required.
The second approach is to use data-driven techniques, which typically leverage (state-of-the-art) SOTA ML model architectures, to address RFF. In the RFF domain, these methods typically employ deep neural networks to extract salient features of individual emitters from the input data, performing automatic RFF whilst simultaneously carrying out a downstream task such as SEI [11]. These DL models have emerged in recent years and have been widely observed delivering superior performance when compared with more traditional methods. Across different types of emitters and signals, these DL approaches are more adaptable for tasks such as signal/modulation-scheme classification, SEI, anomaly detection, amongst others, and have robustness to adverse conditions (e.g. propagation channel and interference) when compared with conventional SP-based approaches [21].
ML, more so DL-based, RFF has the following advantages when compared to conventional SP approaches: (1) It automatically learns nuanced RF fingerprints directly from received transmissions (e.g., from In-phase/Quadrature (I/Q) data streams) whilst simultaneously performing one or more of several possible downstream problems, such as SEI. If SEI is applied, then larger emitter classes can be easily and naturally derived from the detailed RFFs. This circumvents the labour-intensive step of pre-defining them and devising bespoke algorithms to extract them. (2) It enables generic and flexible, multi-purpose, RFF and downstream tasking (RFF-DT), such as SEI, that can be applied to various emitters and signal types including new, previously unseen emitters (with a suitably developed end-to-end ML pipeline). For example, the SEI model can be re-trained/adapted considering the available data from the previously unseen emitters. ML-based RFF-DT, therefore, not only offers a faster development and iterative testing-refinement cycle, but also superior performance (e.g. more accurate identification), including at lower Signal-to-Noise Ratio (SNR) levels. It is emphasised that ML-based RFF strictly refers to automatically extracting the latent RF fingerprint from the input data. This is fundamentally distinct from methods that only apply a ML classifier to an extracted set of features from the received signal (e.g. emitter/signal identification).
The above benefits of ML-based approaches come at the cost of needing to have data (RF captures), with a certain level of data labelling [22, 65], pertaining to the targeted emitters. The quantity of data can also vary depending on the application, but few-shot methods [59] can address this problem should large quantities of data be available for similar types of emitters to those under investigation. In addition to this problem, it is not always guaranteed that the trained model has captured the correct features despite being successful at the task it has been trained to do. For example, the model could latch onto noise profiles in the data which allow it to be successful at the task. Therefore, it is important to ensure that appropriate guardrails are put in place so that this scenario is mitigated.
A block diagram of ML-based RFF-DT is depicted in Figure 1.
@@
The block diagram shows the representation learning module before the downstream task, which is where RF fingerprints are learnt directly from the input data. These fingerprints can have different “levels of supervisions”, referring to the relationship of the RF fingerprint representation learning block and the downstream task block. Most importantly, a general model can be learnt for RFF across numerous emitters or conditions, which can then be adapted or fine-tuned to a specific set of emitters or conditions. The downstream task block can be one or more of the following tasks: (1) Identification of emitters, inclusive of SEI; (2) Comparing and matching multiple emitter emissions originating from the same emitter (e.g. EDA, typically solved using contrastive learning [7]); (3) Automatic grouping and building of RF fingerprint libraries for unknown emitters [27, 53]; (4) Anomaly detection based on learning RFF “normalcy” models, or a generative approach (e.g. variational auto encoders (VAEs)) for detecting and spoofing [70]; (5) Authentication (physical-layer) by comparing RF fingerprints to received signals [37]. In this paper, the ML-based RFF problem is mathematically framed. The main contributions in this paper are as follows:
@@
    A generic framework for ML-based RFF is introduced, providing a basis for treating existing or future novel RFF-related downstream tasks.
@@
    Individual RFF downstream tasks are formulated within the presented framework and solved using a variety of real world datasets, including Automatic Identification System (AIS) data from spaceborne receivers, Digital Mobile Radio (DMR) data and drone RF links data.
@@
    ML-based RFF model performance is demonstrated on real RF data for a variety of tasks and model architectures, spanning across a number of use cases including Spaceborne Intelligence Surveillance and Reconnaissance (ISR), Signals Intelligence (SIGINT) and Counter Unmanned Aerial Systems (C-UAS).
  4. 4.
    Review of existing SOTA RFF ML-based methods, with a focus on input representations, challenges in closed set and open set downstream tasks, as well as mitigations towards the impact of the propagation channel and responses of various receivers.
2 Overview and Related Work
2.1 ML-based RFF approaches
A large corpus of research on ML-based RFF exists in the general wireless communications community, especially for Internet of Things (IoT) and cyber security applications [62]. Various DL models have been adopted and tailored for RFF. For example, convolutional neural networks (CNNs) [66], recurrent neural networks (RNNs) [38], long short-term memory (LSTMs) [50], transformers [55], generative adversarial networks (GANs) [39], auto-encoders (AEs) [67], variational auto-encoders (VAEs) [23], amongst many others have all been applied towards RFF-based problems. CNNs in particular, including complex valued CNNs (CV-CNNs) [24], are amongst the most widely used model architectures. This is generally attributed to their wide availability across different ML libraries/tools and ability to capture temporal correlations in the processed RF signals.
2.2 Closed-set vs Open-set Tasks
The various RFF-based downstream tasks (with the exception of unsupervised learning approaches) can be viewed in terms of being a:
  1. 1.
    Closed-set Downstream Task (CS-DT) : constrained to only consider classes that are within a predefined set of emitters, for which there is existing labelled data/recordings;
  2. 2.
    Open-set Downstream Task (OS-DT) : new, previously unseen during training, emitters are accounted for. These methods recognise and address them appropriately.
2.3 Input Representations
Broadly speaking, ML-based RFF-DT methods can be categorised by their input representation:
  1. 1.
    Time-domain signal: refers to the received I/Q signals of the detected transmission from the targeted emitter and/or their components from empirical mode decomposition, variational mode decomposition, etc. CNNs and RNNs are a common choice for these types of RFF algorithms, including CV-CNNs and LSTMs. Directly using the I/Q data of the detected transmission is highly desirable for a multi-purpose RFF capability for various emitters and scenarios.
  2. 2.
    Spectrogram or transform-domain: bi-spectrum analysis [32] and short-time Fourier Transformer (STFT) [51], Hilbert-Huang [33], and wavelet transforms [63] as well as histograms and grouping input data fall into this category. The resultant representations (or information extracted from them) are then fed into a neural network, often a CNN. Any such transformation should preserve the relevant information in the (complex) time domain RF data to faciliate the superior feature extraction capability of the DL algorithms.
  3. 3.
@@ 2.3 Input Representations
    Constellation-based: relies on the differential constellation trace figure, often based on the modulation scheme adopted by the emitter [42]. This modality is applicable to communication signals and these constellation traces typically serve as an input to a multi-channel CNN. However, emitter classification from the constructed constellation trace can be challenging in practice and will be specific to an emitter type/modulation scheme, i.e., leads to an inflexible RFF capability.
Time series inputs can include either the transient or the steady-state, or both, components of the received signal. The transient is related to when the transmitter turns on (e.g. beginning of the received burst) and usually incorporates more prominent unique intrinsic distinguishing features originating from hardware imperfections [54]. The steady-state segment often contains the communication content, if relevant, and follows from the transient phase. As well as message-related consistent patterns (e.g. preamble/training-sequence, if relevant), steady-state segments still contain salient transmitter features that can be useful.
2.4 Mitigating Impact of Propagation Channel and Responses of Various Receivers
@@
In practice, the received signal can be substantially impacted by multipath, fading and noise from the propagation channel. Similarly, inherent differences amongst various deployed receivers (e.g. distributed on different satellites in a constellation) can induce a significant mismatch (e.g. distribution shift) between the training dataset and the signals that the ML RFF approach must handle when deployed. Naturally, these scenarios degrade the performance of any RFF-based model. Therefore, successful RFF-based models need to incorporate, where needed, model adaptation, domain adaptation, transfer learning and few-shot learning approaches. All of these approaches enable generalisable and robust ML-based RFF as well as SEI [68, 29, 58, 60]. The aim of these approaches is to improve the resilience of the model in light of certain environmental and sensing conditions. These methods have been tested and applied in numerous research papers on real and synthetic datasets that pertain to complex environments (e.g. dense urban or indoor settings) and for transmission frequencies which are particularly susceptible to channel effects (e.g. deep fading such as with ZigBee [66], Bluetooth low energy [3] and Wi-Fi [17]).
3 Problem Formulation

### v2

@@ Abstract
Fingerprinting radio frequency (RF) emitters typically involves finding unique characteristics that are featured in their received signal. These fingerprints are nuanced, but sufficiently detailed, motivating the pursuit of methods that can successfully extract them. The downstream task that requires the most meticulous RF fingerprinting (RFF) is known as specific emitter identification (SEI), which entails recognising each individual transmitter. RFF and SEI have a long history, with numerous defence and civilian applications such as signal intelligence, electronic surveillance, physical-layer authentication of wireless devices, to name a few. In recent years, data-driven RFF approaches have become popular due to their ability to automatically learn intricate fingerprints. They generally deliver superior performance when compared to traditional RFF techniques that are often labour-intensive, inflexible, and only applicable to a particular emitter type or transmission scheme. In this paper, we present a generic and versatile machine learning (ML) framework for data-driven RFF with several popular downstream tasks such as SEI, data association (EDA) and RF emitter clustering (RFEC). It is emitter-type agnostic. We then demonstrate the introduced framework for several tasks using real RF datasets for spaceborne surveillance, signal intelligence and countering drones applications.
I Introduction
I-A Background and Overview
@@
RFF involves building representations or profiles of RF emitters from distinct features that are present in their received transmissions. These can be utilised by a downstream task such as identifying each individual transmitter or distinguishing between emitter types (e.g. different radars, 5G sources, etc.) [1, 2]. In general, the sought emitters’ characteristics can be intentional (e.g. the waveform parameters, modulation scheme, transmission timing or bandwidth, etc.), albeit impacted by the propagation channel effects (e.g. fading, dispersion, scintillation, etc.), or unintentional [1, 3]. The latter often originate from inherent transmitter hardware imperfections (e.g. amplitude rise-time, ripple, phase distortion, nonlinearities in the processing chain, antenna responses, etc.) and/or are induced by manufacturing variabilities, including from individual hardware components [1, 4, 5, 2]. Spoofing or obfuscating the resultant (physical-layer) fingerprints is difficult. Overall, RFF entails defining and extracting salient features (intentional or not) that enable the recognition of the targeted emitters [6, 4, 2, 7], this is, RF fingerprint identification (RFFI).
In practice, data captures from a wide range of modulation schemes and transmission protocols are collected for coarse RFFI problems such as conventional automatic modulation or protocol identification [8, 9]. The primary focus in this paper is fingerprinting that permits solving more granular recognition tasks such as specific emitter identification. This is because the fingerprints in this case need to incorporate the most comprehensive and nuanced set of characteristics attributed to each individual emitter, typically exploiting unintentional features originating from hardware imperfections and/or manufacturing variabilities [1, 4, 2]. For example, different instances of the same transmitter (i.e. same device model, make, transmission protocol, modulation scheme, etc.) can be uniquely identified with SEI; see the digital mobile radio and drones datasets in Section V. Note that broader emitter classes (e.g. individual manufacturers) and even larger categories (e.g. modulation or protocol used), can be naturally derived and/or inform, implicitly or explicitly, by the detailed RFF for SEI.
In general, there are two broad categories of RFF methods:
@@
    Traditional RFF with feature engineering and signal processing extraction techniques [10, 11, 12, 13, 14, 15].
@@
    Data-driven, ML, typically deep learning (DL)-based, RF fingerprinting [16, 17, 9, 18, 19, 20, 21, 22, 23].
Traditional RFF relies on manually defining the distinguishable features/characteristics that form the emitters’ RF fingerprints such as evolution of the signal’s frequency spectrum, central frequency offsets or its amplitude envelope [1, 20, 22, 4]. Subsequently, a suitable signal processing (SP) algorithm is applied to obtain them from received transmissions. The fingerprint features are then ingested by a conventional classifier (inclusive of rule-based or classical machine learning methods such as support vector machines, decision trees, multi-layer perceptrons, etc.) for identification.
@@
In general, traditional RFF methods can be effective, computationally efficient and their outputs fully explainable. However, they are inherently inflexible and usually specific to each emitter/signal type and a particular RFFI problem. They are labour-intensive requiring considerable RF specialists’ effort to manually define the signal characteristics that constitute the RF fingerprint of each new type of emitter/signal and involve adapting/devising tailored SP techniques. This renders their development cycle slow and expensive. This can be problematic in the increasingly congested and contested electromagnetic environment (EME), where new emitters emerge regularly and communication technologies continuously evolve. Therefore, a more versatile and flexible approach to RFF might be desired.
On the other hand, data-driven RFF techniques leverage state-of-the-art (SOTA) ML, typically with deep neural networks (DNNs). They have emerged in recent years as a promising RFF methodology due to the superior performance (especially SEI) they offer [24, 25, 1, 20, 21, 22, 4]. DL-based RFF advantages are threefold. First, it automatically learns intricate fingerprint features from the data, for instance from in-phase/quadrature (I/Q) signals, whilst simultaneously performing one or more of several possible downstream tasks such as SEI, emitter data association and RF emitter clustering. This circumvents the aforementioned labour-intensive steps of the traditional RFF (i.e. defining fingerprint features and developing bespoke SP algorithms). Second, data-driven fingerprinting enables generic, multi-purpose and emitter-type agnostic, fingerprinting. It can be applied to various emitters and signal types including new, previously unseen emitters (with a suitably developed end-to-end ML pipeline). Third, ML-based RFF provides additional robustness against adverse channel conditions and interference [1, 20, 4]. This facilitates emitter identification at lower signal-to-noise ratios (SNRs) compared to competing approaches, including those that reply on fully or partially demodulating-decoding the received signals from (non-)cooperative transmitters.
A block diagram of ML-based RFF is depicted in Figure 1. The representation learning module is where the RF fingerprints are learnt directly from the input data. This can be carried out under different “levels of supervisions”, referring to the relationship of the RF fingerprint representation learning block and the downstream task block that can be one or more of the following: (1) Identification of emitters, inclusive of SEI; (2) Comparing and matching multiple emitter emissions originating from the same emitter (e.g. EDA, typically solved using contrastive learning [26]); (3) Automatic grouping and building of RF fingerprint libraries for unknown emitters [27, 28]; (4) Anomaly detection based on learning RFF “normalcy” models, or a generative approach, e.g. variational auto-encoders (VAEs) [29]; (5) Authentication (physical-layer) by comparing known fingerprints from those from received signals [30]. Most importantly, a general RFF model can be learnt across numerous emitters or conditions, which can then be adapted or fine-tuned to a specific set of scenarios. This RFF downstream tasking (RRF-DT) has several applications such as for Signal Intelligence SIGINT (e.g. to detect-track threats), cyber security for wireless devices and networks (e.g. for internet of things IoT), Intelligence, Surveillance and Reconnaissance (ISR) and others [1, 3, 4, 2, 20].
ML-based RFF-DT, therefore, not only offers a faster development and iterative testing-refinement cycle, but also SOTA performance.  This comes at the cost of needing to have access to sufficient representative data from the targeted emitters, with a certain level of ground-truth labels (i.e. supervision, if any) dependent on the the downstream task(s) [31, 32]. Nevertheless, few-shot learning methods permit adapting a model trained on data from similar types of emitter(s) to those under investigation from a modestly small number of example RF captures [20, 1, 4]. As is common in adopting data-driven, DL, methods there is a risk that the model does not learn the sought RF fingerprint (e.g. hardware-imperfections-related for SEI). Instead, it could latch onto the noise or EME profile or transmission pattern (e.g. frequency hopping) in the available data. To ensure mitigating such overfitting effects, it is crucial to employ appropriate best practices in ML/DL [33] such as cross validation, model simplification, data augmentation with synthetic examples, loss function regularisation, early stopping criteria for optimisers, ensemble methods, etc.
@@
I-B Contributions
The main contributions of this paper are:
  • A generic (mathematical) framework for ML-based RFF is introduced, providing a basis for treating various downstream tasks.
  • A number of key individual RFF-related tasks (namely SEI, EDA and RFEC) are then formulated within this framework and solved using a variety of real RF datasets from automatic identification system (AIS) as captured by spaceborne receivers, digital mobile radios (DMRs) and drone RF links.
  • Performance of the ML-based RFF-DT is demonstrated with several model architectures; this is in the context of a number of use cases such as for ISR, SIGINT and counter unmanned aerial systems (C-UAS).
This is preceded by a brief overview of SOTA ML-based RFF, encompassing input representations, challenges in closed set and open set problems, as well as mitigating the impacts of channel conditions and/or receivers responses.
I-C Paper Layout
The remainder of the paper is organised as follows. In Section II, a concise summary of ML-based RFF and associated challenges are presented. Whilst Section III outlines the overall RFF framework, individual downstream tasks are stated in Section IV. The various use cases and real RF datasets are described in Section V and results are discussed in Section VI. Finally, conclusions are drawn in Section VII.
II Related Work
A large corpus of research on ML-based RFF exists in the general wireless communications community, especially for IoT and cyber security applications [6, 34, 20, 7]. Various DL models have been adopted and tailored for RF fingerprinting and identification. For example, convolutional neural networks (CNNs) [35], recurrent neural networks (RNNs) [36], long short-term memory (LSTMs) [18], transformers [37], generative adversarial networks (GANs) [38], auto-encoders (AEs) [39], variational auto-encoders (VAEs) [40], amongst many others have all been applied towards RFF-based problems. CNNs in particular, including complex valued CNNs (CV-CNNs) [41], are amongst the most widely used model architectures. This is generally attributed to their wide availability across different ML libraries/tools and ability to capture temporal correlations in the processed RF signals. Next, we highlight a few key aspects associated with RFF inclusive of the nature of the downstream problem, input representations and impact of channel effects.
II-A Closed-set vs Open-set Tasks
The various RFF-based downstream tasks, except those that use unsupervised learning, can be viewed as being a:
@@ II-A Closed-set vs Open-set Tasks
    Closed-set downstream task (CS-DT): constrained to a predefined finite set of emitters, for which there is existing labeled data/recordings.
@@
    Open-set downstream task (OS-DT): new, previously unseen (i.e. in model training), RF emitters can be encountered, requiring suitable techniques to recognise their presence and address them appropriately (e.g. identify them if detected again, if relevant).
II-B Input Representations
Broadly speaking, ML-based RFF-DT methods can be categorised by their input:
  1. 1.
    Time-domain signal: refers to the received I/Q data of the detected transmissions from the targeted emitter and/or their resultant components from empirical mode decomposition, variational mode decomposition, etc. CNNs and RNNs are a common choice for this data representation, including CV-CNNs and LSTMs. Directly ingesting the I/Q data can be highly desirable for a multi-purpose, versatile, RFF capability for various emitter types and scenarios.
  2. 2.
    Spectrogram or transform-domain: bi-spectrum analysis [46] and short-time Fourier transformer (STFT) [47], Hilbert-Huang [48], and wavelet transforms [49] as well as histograms and data groupings fall into this class. The transform representations (or information extracted from them) are then fed into a neural network, e.g. a CNN. Any transformation should preserve the relevant information in the (complex) time domain RF data to facilitate the pursued feature extraction capability of the DL algorithms.
@@ II-B Input Representations
    Constellation-based: relies on the differential constellation trace figure, often based on the modulation scheme employed by the emitter [9]. This modality is applicable to communication signals and the constellation traces typically serve as an input to the DL model (e.g. a multi-channel CNN). However, emitter classification from the constructed trace can be challenging in practice and is specific to an emitter type/modulation-scheme, i.e., leads to an inflexible RFF capability.
The time series inputs can include the transient and/or the steady-state segments of the received signal [6, 20, 22]. The former is related to the start of the transmission (e.g. initial part of a burst inclusive of its rising edge) and often incorporates prominent unique distinguishable characteristics related to the emitter’s hardware. Accurate detection of the start of emissions can be difficult, especially in challenging EMEs. The steady-state segment follows from the transient phase and can include message/payload-related data, if relevant. It still nonetheless contains salient emitter features that enable SEI [20, 22]. Processing a combination of transient and steady-state components of the transmission can often be the pragmatic approach in practice.
II-C Mitigating Channel and Receiver Response Effects
@@
In practice, the received signal can be substantially impacted by multipath, fading and noise from the propagation channel. Similarly, inherent differences amongst various deployed receivers (e.g. distributed on different satellites in a constellation) can induce a significant mismatch (e.g. distribution shift) between the training dataset and the signals that the ML RFF approach must handle when deployed. Naturally, these scenarios degrade the performance of any RFF-based model. Therefore, successful RFF-based models need to incorporate, where needed, model adaptation, domain adaptation, transfer learning and few-shot learning approaches. All of these approaches enable generalisable and robust ML-based RFF as well as SEI [50, 51, 52, 53]. The aim of these approaches is to improve the resilience of the model in light of certain environmental and sensing conditions. These methods have been tested and applied in numerous research papers on real and synthetic datasets that pertain to complex environments (e.g. dense urban or indoor settings) and for transmission frequencies which are particularly susceptible to channel effects, for example deep fading such as with ZigBee [35], Bluetooth low energy [54] and Wi-Fi [55].
III Problem Statement and RFF Framework